代数解析の未来 (Special Project) organised by T. Tanisaki and K. Takeuchi Date: １月１５日 (火) – １月１８日 (金) , 2002 Place: 京都大学数理解析研究所, 202 号室 (2F)

Jan 15 (Tuesday) 13:25〜13:30 Opening ceremony 13:30〜14:20 P. Schapira (Paris 6, France) Review on ind-sheaves （with M. Kashiwara） 14:30〜15:20 K. Vilonen （Northwestern univ, USA） Character formulas for representations of real Lie groups (with W. Schmid) 15:30〜16:20 Mutsumi Saito （Hokkaido univ） Isomorphism classes of A-hypergeometric systems

Jan 16 (Wednesday) 10:00〜10:50 P. Schapira (Paris 6, France) Micro-support and regularity for ind-sheaves, applications to D-modules and open problems （with M. Kashiwara） 11:00〜11:50 N. Takahashi (Hiroshima univ.) An algebra of differential operators on the formal superloop space 13:30〜14:20 T. Kakehi (Univ. of Tsukuba) Some estimates for Radon transforms 14:30〜15:20 A. D’Agnolo (Paris 6, France) Radon and Fourier transform for D-modules 15:30〜16:20 C. Marastoni （Padova univ, Italy） Duality for generalized flag manifolds and D-modules

Jan 17 (Thursday) 10:00〜10:50 K. Vilonen （Northwestern univ, USA） Geometric Satake isomorphism (with I. Mirkovic) 11:00〜11:50 S. Tajima (Niigata univ.) and Y. Nakamura (Ochanomizu univ.) Toward the computation of the Grothendieck local residues 13:30〜14:20 K. Takeuchi (Univ. of Tsukuba) On a certain class of holonomic D-modules 14:30〜15:20 Y. Sugiki (Univ. of Tokyo) The dual concept of sheaves and Integral transfroms 18:00〜 reception Jan 18 (Friday) 10:00〜10:50 A. Kamita (Hiroshima univ.) Generalized universal Verma modules and b-functions 11:00〜11:50 T. Oaku (Tokyo woman’s Christian univ.) and N. Takayama (Kobe univ.) Minimal filtered free resolutions of homogenized D-modules 13:30〜14:20 J-P. Schneiders （Paris 13, France） Sheaves with values in quasi-abelian categories 14:30〜15:20 J-P. Schneiders （Paris 13, France） Topological sheaves and reconstruction theorem for D^{∞}-modules (with F. Prosmans) 15:30〜16:20 K. Vilonen （Northwestern univ, USA） On the geometric Langlands conjecture